OpenStudy (anonymous):

i have a non-homogeneous equation.. And I need to take a guess for Yp. The r(x) = x*exp(3x), i'm not sure what guess for Yp to take

OpenStudy (anonymous):

what is Yp

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx If you look at example one, a guess for Yp is taken based on thier example The actual equation for me is y'' -4y' + 4y = xexp(3x) So far I have Yh = Aexp(2x) + Bexp(2x) Now I need to find Yp..

OpenStudy (anonymous):

woops, Yh = Aexp(2x) + Bxexp(2x)

OpenStudy (anonymous):

i m puzzled...........

OpenStudy (anonymous):

let yp=xe^3x(ax+b) now differentiating y''=e^3x(9ax+9b+6a) y'=e^3x(3ax+3b+a) and putting the values in y''-4y'+4y=xe^3x you get e^3x((9ax+9b+6a)-(12ax+12b+4a)+(4ax+4b))=xe^3x ax+b+2a=x so a=1, b+2a=0 => b=-2 yp=e^3x(x-2)